Discuss the continuity of f(x, y) = (x^4)(y^4)/(x2+y 2)^3 if (x, y) not equal to (0,0), 0 if (x, y) = (0,0) at(0, 0) 


1
Expert's answer
2021-11-08T08:19:13-0500

Let "x=r\cos \theta, y=r\sin \theta." Then


"\dfrac{x^4y^4}{(x^2+y^2)^3}=\dfrac{(r^4\cos^4\theta) (r^4\sin^4\theta)}{(r^2\cos^2\theta+r^2\cos^2\theta)^3}="

"=r^2\cos^4\theta\sin^4\theta"

"\lim\limits_{(x,y)\to(0,0)}\dfrac{x^4y^4}{(x^2+y^2)^3}=\lim\limits_{r\to0}(r^2\cos^4\theta\sin^4\theta)=0"

We have that


"\lim\limits_{(x,y)\to(0,0)}f(x,y)=0=f(0,0)"

Therefore the function


"f(x, y)= \begin{cases}
 \dfrac{x^4y^4}{(x^2+y^2)^3} &\text{if } (x,y)\not=(0,0) \\
 0 &\text{if } (x,y)=(0,0)
\end{cases}"

is continuous at "(0,0)."